OpenStudy (anonymous):

Help with Questions Please!! Please do NOT tell me the Answer Just tell me How to solve! Thank you. 1. Evaluate the function y = -2x + 5 for x values of -1,0, 1,2, and 3. 2. You deposit $2,500 in an account that pays 5.7% interest compounded anually. What is the balance after five years? Solve each formula for x. 4. z = 3x + y 5. k = -4xyz 6. 1/9x - 4 = z/3

OpenStudy (anonymous):

1. substitute values of x given in the equation and calculate y 2. Calculate how much is 5,7% of 2500. Add it to 2500. Calculate 5,7% of the result and add it. Do this 5 times 3. add costs and benefits and devide by amount of people 4. transform each equation so that on the left side of the equality you have only x and at the right side all the rest

OpenStudy (diyadiya):

@myko is right!

OpenStudy (anonymous):

other questions?

OpenStudy (anonymous):

5,6?

OpenStudy (diyadiya):

5 and 6 to be done in the same way as 4

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Can you do it?

OpenStudy (anonymous):

I'm stuck on #4-6

OpenStudy (diyadiya):

4. z = 3x + y subtract y from both sides z-y=3x+y-y z-y=3x divide by 3 (z-y)/3=x x=(z-y)/3

OpenStudy (anonymous):

How about now?

OpenStudy (anonymous):

yea i think i get it know